OpenStudy (anonymous):

How you turn this 1/2PI -1/2[sin(2+PI)-sin2]=1/2PI+sin2? I used the property of sinx-siny=2[sin((x-y)/2)][cos((x+y))].. But i cant get the answer

OpenStudy (anonymous):

Let me clarify, you want to prove the trigonometric identity?

OpenStudy (anonymous):

oh wait nvm that is incredibly simple

OpenStudy (anonymous):

ok so take a look at sin(2-PI), Your goal is to rewrite that so you can reduce the equation. Using CAST rule, you will see that adding/subtracting PI will make the sine negative Therefore: sin(2-PI) = -sin(2) Plug that back in and reduce the terms, you get the exact equation. Do you want a diagram?

OpenStudy (anonymous):

1/2PI -1/2[sin(2+PI)-sin2]=1/2PI+sin2 1/2PI -1/2[-sin(2)-sin2]=1/2PI+sin2 etcetera

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

sin(2-pi) = -sin(2).. i don understand this part, there is no identify of such "?

OpenStudy (anonymous):

and thanks alot for the explanation (:

OpenStudy (anonymous):

can u explain, how u get sin(2-pi) = -sin(2)? actually that is the only thing that confused me in this transformation..

OpenStudy (amistre64):